A 0.350 m gold bar at 20.0°C heats up to 500.0°C. What is the change in length of the gold bar (
1
Expert's answer
2020-11-27T10:08:42-0500

We can find the change in length of the gold bar from the formula:


"\Delta L=\alpha L\Delta T,""\Delta L=14\cdot10^{-6}\ ^{\circ}C^{-1} \cdot 0.350\ m\cdot 480\ ^{\circ}C=2.35\cdot10^{-3}\ m."

Answer:

"\Delta L=2.35\cdot10^{-3}\ m."